### v1

# Chern character of Schur bundles

Abstract.
Given a vector bundle $E$, we give an explicit formula to compute Chern classes of Schur bundles $\mathsf{\mathbf{S}}^{\alpha}E$ in terms of those of $E$.
Introduction
Given a complex holomorphic vector bundle $E$ on a compact complex manifold $X$, one can associate to it a Schur bundle
$\mathsf{\mathbf{S}}^{\alpha}E,$
where $\alpha$ is a Young diagram of length at most $\operatorname{rk}(E)$. If we denote by $V$ a fibre of $E$, then a fibre of $\mathsf{\mathbf{S}}^{\alpha}E$ is the irreducible Schur $\operatorname{GL}(V)$-module $\mathsf{\mathbf{S}}^{\alpha}V$.
An important problem in computational algebraic geometry is to compute the Chern classes of $\mathsf{\mathbf{S}}^{\alpha}E$ in terms of the Chern classes of $E$. It is in principle possible to do such a computation by using the splitting principle, but the combinatorics behind this approach is very cumbersome, see for example [Svr93], where the case of symmetric and alternating powers was treated. In this short paper we provide a simple solution to this problem.
Theorem (Theorem 3.2).
For any $n\geq 0$, the $n^{\operatorname{th}}$-Chern character coefficient of $\mathsf{\mathbf{S}}^{\alpha}E$ is
$\operatorname{ch}_{n}(\mathsf{\mathbf{S}}^{\alpha}E)=\sum_{\nu\vdash n}\operatorname{rk}(\mathsf{D}^{\nu}(\mathsf{\mathbf{s}}_{\alpha}))\operatorname{ch}^{(\nu)}(E).$
Let us explain the formula above:
  • the sum runs over all partitions $\nu=(1^{\nu_{1}},2^{\nu_{2}},\dots)$ of the number $n$;
  • $\mathsf{\mathbf{s}}_{\alpha}$ is the Schur polynomial in $\operatorname{rk}(E)$ variables associated to the Young diagram $\alpha$;
  • $\mathsf{D}^{\nu}(\mathsf{\mathbf{s}}_{\alpha})$ is the $\nu$-derived symmetric polynomial which we define in Section 2.3;
  • $\operatorname{rk}\colon\mathbb{Q}[x_{1},\dots,x_{r}]^{\mathfrak{S}_{r}}\to\mathbb{Q}$ is the function associating to each Schur polynomial the dimension of the corresponding irreducible polynomial $\operatorname{GL}(r)$-representation;
  • $\operatorname{ch}^{(\mu)}(E)=\frac{\operatorname{ch}_{1}(E)^{\mu_{1}}}{\mu_{1}!}\frac{\operatorname{ch}_{2}(E)^{\mu_{2}}}{\mu_{2}!}\cdots$ are the divided powers of the Chern character coefficients.
The quantity $\mathsf{D}^{\nu}(\mathsf{\mathbf{s}}_{\alpha})$, the new input in the formula, is a purely combinatorial object which is easily computed in explicit examples, see Examples 3.5 and 3.6 and Proposition 3.8.
The proof of Theorem 3.2 is simple and only uses some weak properties of the Chern character, like multiplicativity and additivity. To emphasise this aspect, and thus making the proof more transparent, we undress it of all the superficial structures and reduce it to a statement about polynomials in countably many variables. This is the content of our Key Lemma (Lemma 1.3), from which all the results in the paper are deduced.
Notice that this abstract point of view is not only an exercise of style. The same proof in fact ensures that the result holds in any category where suitable characteristic classes can be defined. For example, it holds for proper algebraic varieties and locally free sheaves over them. Similarly, it also holds for characteristic classes valued in a Chow ring, instead of a cohomology ring. As an example of characteristic classes that differs from the Chern character, we also prove the analog statement for the twisted Chern character $k(F)=\exp\left(-\frac{c_{1}(F)}{\operatorname{rk}(F)}\right)\operatorname{ch}(F)$ considered in [Mar20] as characteristic classes of projective bundles (see Theorem 4.1).
As already mentioned, for symmetric and exterior powers a concrete formula was found by Svrtan in [Svr93]. We give alternative formulas for these special Schur bundles in Proposition 3.8 (see Remark 3.11). The formula for the first Chern class of a Schur bundle was a folklore result, whose first rigorous proof can be found in [Rub13]. We present in Corollary 3.3 an alternative proof.
In our previous work [DFO25a] (see also [DFO25b, Appendix A]) we computed the second and third Chern characters of Schur functors in a similar spirit to the one presented here. This work extends and complements those results, see § 4.
Plan of the paper
The technical core of the paper is contained in § 1, where the Key Lemma is proved. In § 2 we recover some important facts about the ring of symmetric functions, that will be used in § 3, where the Theorem is proved. In § 4 we make explicit the example of the twisted Chern character.
Acknowledgements
This research has been partially funded by the European Union - NextGenerationEU under the National Recovery and Resilience Plan (PNRR) - Mission 4 Education and research - Component 2 From research to business - Investment 1.1 Notice Prin 2022 - DD N. 104 del 2/2/2022, from title “Symplectic varieties: their interplay with Fano manifolds and derived categories”, proposal code 2022PEKYBJ – CUP J53D23003840006. The second and third authors are members of the INDAM-GNSAGA group.
1. The Key Lemma
Let $R=\mathbb{Q}[x_{1},x_{2},\dots]$ be a polynomial ring in countably many variables $x_{k}$, $k\geq 1$; we think of $R=\bigoplus_{d\geq 0}R_{d}$ as a graded ring by setting the degree of each indeterminate to be $1$. We are also given a commutative unital graded $\mathbb{Q}$-algebra $A=\bigoplus_{n\geq 0}A_{n}$, where $\mathbb{Q}\subset A_{0}$, whose completion we denote by $\widehat{A}=\prod_{n\geq 0}A_{n}$. Finally, we fix a collection of positive rational numbers $\mathsf{q}=\{q_{n,k}\}_{n\geq 1,\,k\geq 1}$ with $q_{n,1}=1$.
Definition 1.1.
Consider a map $\mathsf{C}:R\to\widehat{A}$ whose projection to each graded component we denote by $\mathsf{C}_{n}\colon R\to A_{n}$, so that
$\mathsf{C}=\sum_{n\geq 0}\mathsf{C}_{n}.$
We say that $\mathsf{C}$ is an $\widehat{A}-$valued $\mathsf{q}$-characteristic class if
  1. (1)
    $\mathsf{C}_{0}$ is a ring homomorphism;
  2. (2)
    $\mathsf{C}$ is multiplicative, i.e. $\mathsf{C}(fg)=\mathsf{C}(f)\mathsf{C}(g)$ for all $f,g\in\mathbb{Q}[x_{1},x_{2},\dots]$;
  3. (3)
    the restriction of $\mathsf{C}$ to each graded component $R_{d}$ is $\mathbb{Q}$-linear;
  4. (4)
    for every $n\geq 1$ and every $k\geq 1$, we have
    $\mathsf{C}_{n}(x_{k})=q_{n,k}\mathsf{C}_{n}(x_{1}).$
Remark 1.2.
We refer to the property in (2) above as same-degree additivity. For our main application, $\mathsf{C}$ will be the Chern character, which is additive; nevertheless, allowing this more general setting gives us the possibility to extend the results to other classes, see for example § 4.
For each $n\geq 0$, let us now consider the derivation
$\mathsf{D}_{n}=\sum_{k\geq 1}q_{n,k}\frac{\partial}{\partial x_{k}}\in\operatorname{Der}R.$ (1.1)
Notice that $\mathsf{D}_{n}$ is locally nilpotent and well-defined, since for each polynomial $f\in\mathbb{Q}[x_{1},x_{2},\dots]$ only finitely many of the $\frac{\partial f}{\partial x_{k}}$ are non-zero.
The formal sum $\sum_{n\geq 1}\mathsf{C}_{n}(x_{1})\cdot\mathsf{D}_{n}$ is then a (locally nilpotent, converging) derivation of $R_{\widehat{A}}=\widehat{A}\otimes_{\mathbb{Q}}R=\widehat{A}[x_{1},x_{2},\dots]$ and its exponential is a well defined automorphism of $R_{\widehat{A}}$, whose restriction to $R\subset R_{\widehat{A}}$ yields a $\mathbb{Q}$-algebra homomorphism
$\exp\left(\sum_{n\geq 1}\mathsf{C}_{n}(x_{1})\cdot\mathsf{D}_{n}\right)\colon\mathbb{Q}[x_{1},x_{2},\dots]\longrightarrow R_{\widehat{A}}=\widehat{A}\otimes_{\mathbb{Q}}\mathbb{Q}[x_{1},x_{2},\dots].$
Finally, we abuse notation and also denote by $\mathsf{C}_{0}$ the composition
$\widehat{A}\otimes_{\mathbb{Q}}\mathbb{Q}[x_{1},x_{2},\dots]\stackrel{{\scriptstyle 1\otimes\mathsf{C}_{0}}}{{\longrightarrow}}\widehat{A}\otimes_{\mathbb{Q}}A_{0}\stackrel{{\scriptstyle\mu}}{{\longrightarrow}}\widehat{A}$,
where $\mu$ is the multiplication map of $\widehat{A}$. Notice that since $\widehat{A}$ is commutative, the multiplication map $\mu$ is a ring homomorphism; consequently, the composition
$\mathsf{C}_{0}\circ\exp\left(\sum_{n\geq 1}\mathsf{C}_{n}(x_{1})\cdot\mathsf{D}_{n}\right)\colon R\longrightarrow\widehat{A}$
is a well-defined $\mathbb{Q}$-algebra homomorphism by construction.
Lemma 1.3 (Key Lemma).
Let $\mathsf{C}$ be a $\widehat{A}-$valued $\mathsf{q}$-characteristic class. Then,
$\mathsf{C}=\mathsf{C}_{0}\circ\exp\left(\sum_{n\geq 1}\mathsf{C}_{n}(x_{1})\cdot\mathsf{D}_{n}\right),$ (1.2)
when both sides are applied to any given homogeneous polynomial in $R$. If furthermore $\mathsf{C}$ is a $\mathbb{Q}$-algebra homomorphism, then equality (1.2) holds in general.
Proof.
Let us prove that (1.2) holds when applied to the element $x_{k}$. The left-hand side is
$\mathsf{C}(x_{k})=\sum_{n\geq 0}\mathsf{C}_{n}(x_{k})=\mathsf{C}_{0}(x_{k})+\sum_{n\geq 1}q_{n,k}\mathsf{C}_{n}(x_{1}).$
As for the right-hand side, we notice that $\mathsf{D}_{n}(x_{k})=q_{n,k}\in\mathbb{Q}$, so that $\mathsf{D}_{m}(\mathsf{D}_{n}(x_{k}))=0$ for all $m$ and $n$. When we apply the automorphism to $x_{k}$ only the constant and the linear term survive and we obtain
$\exp\left(\sum_{n\geq 1}\mathsf{C}_{n}(x_{1})\cdot\mathsf{D}_{n}\right)(x_{k})=x_{k}+\sum_{n\geq 1}q_{n,k}\mathsf{C}_{n}(x_{1}).$
After applying $\mathsf{C}_{0}$, this coincides with the left-hand side.
By multiplicativity of $\mathsf{C}$, Equation (1.2) holds for each monomial, and since the restriction of $\mathsf{C}$ on each graded component of $R$ is $\mathbb{Q}$-linear, it also holds for every homogeneous polynomial. ∎
For the following corollary we use the notation $\nu=(1^{\nu_{1}},2^{\nu_{2}},\dots,h^{\nu_{h}})$ for partitions. In particular, $\nu\vdash n$ means that $\sum_{i}i\nu_{i}=n$.
If $\nu=(1^{\nu_{1}},2^{\nu_{2}},\dots,h^{\nu_{h}})$ is a partition, then we put
$\mathsf{D}^{\nu}=D_{1}^{\nu_{1}}\circ\cdots\circ\mathsf{D}_{h}^{\nu_{h}}.$ (1.3)
Notice that, for any $n,m\in\mathbb{Z}$, the operators $\mathsf{D}_{n}$ and $\mathsf{D}_{m}$ commute, i.e. $\mathsf{D}_{n}\mathsf{D}_{m}=\mathsf{D}_{m}\mathsf{D}_{n}$.
Corollary 1.4.
Let $\mathsf{C}$ be a $\hat{A}-$valued $\mathsf{q}$-characteristic class. Then for every $n\geq 0$ and every $f\in\mathbb{Q}[x_{1},x_{2},\dots]$
$\mathsf{C}_{n}(f)=\sum_{\nu\vdash n}\mathsf{C}_{0}(\mathsf{D}^{\nu}(f))\mathsf{C}^{(\nu)}(x_{1}),$
where $\mathsf{C}^{(\nu)}=\prod_{i}\frac{\mathsf{C}_{i}^{\nu_{i}}}{\nu_{i}!}$ is the divided power of $\mathsf{C}$. ∎
2. Symmetric functions
In this short section we recall some facts about symmetric functions, see [Mac95] for a reference.
2.1.
The following are the notations we use in this and the next section.
  • Young diagrams will be denoted by the letter $\alpha$. We also represent Young diagrams via vectors whose entries denote the length of rows of $\alpha$. The size $|\alpha|$ of a Young diagram is the number of boxes or, equivalently, the sum of the entries in the vector representation. The length of $\alpha$ is the number of rows or, equivalently, the number of non-zero entries in its vector representation.
  • The notation $(i,j)\in\alpha$ means that we are considering the box in position $(i,j)$ in $\alpha$. The upper-left box has coordinates $(1,1)$ and we increase each coordinate by one at every step downwards or rightwards, accordingly.
  • Given $(i,j)\in\alpha$, we denote by $\operatorname{Hook}_{(i,j)}(\alpha)$ the hook centred in position $(i,j)$ and by $\operatorname{h}_{(i,j)}(\alpha)$ he corresponding hook number, i.e. its size.
  • Partitions will be denoted by $\mu=(1^{\mu_{1}},2^{\mu_{2}},\dots)$ and their size is $|\mu|=\sum_{i}i\mu_{i}$.
  • $\Lambda=\bigoplus_{d}\Lambda^{d}$ is the graded ring of symmetric functions with coefficients in $\mathbb{Q}$.
  • Schur functions are denoted by $\mathsf{\mathbf{s}}_{\alpha}$ and are indexed by Young diagrams. For $|\alpha|=d$ we have $\mathsf{\mathbf{s}}_{\alpha}\in\Lambda^{d}$.
  • Power sums functions are $\mathsf{\mathbf{p}}_{k}(x_{1},x_{2},\dots)=\sum_{i\geq 1}x_{i}^{k}\in\Lambda^{k}$.
  • If $\mu$ is a partition, then we set
    $\mathsf{\mathbf{p}}_{\mu}:=\mathsf{\mathbf{p}}_{1}^{\mu_{1}}\mathsf{\mathbf{p}}_{e}^{\mu_{2}}\cdots\in\Lambda^{|\mu|}$
    for the corresponding monomial in power sums.
$2.$
Let $\Lambda=\bigoplus\Lambda^{d}$ be the graded ring of symmetric functions with coefficients in $\mathbb{Q}$. Two sets of generators of $\Lambda$ are given by Schur functions $\mathsf{\mathbf{s}}_{\alpha}$ and power sums functions $\mathsf{\mathbf{p}}_{k}$. A Schur function $\mathsf{\mathbf{s}}_{\alpha}$ is homogeneous of degree $|\alpha|$, while a power sums function $\mathsf{\mathbf{p}}_{k}$ is homogeneous of degree $k$.
The homogeneous component $\Lambda^{d}$ is generated (as a $\mathbb{Q}$-vector space) by Schur functions $\mathsf{\mathbf{s}}_{\alpha}$ with $|\alpha|=d$ or by monomials in power sums $\mathsf{\mathbf{p}}_{\mu}$ with $|\mu|=d$.
There is a natural isomorphism
$\Lambda\cong\mathbb{Q}[\mathsf{\mathbf{p}}_{1},\mathsf{\mathbf{p}}_{2},\dots]$
exhibiting the ring of symmetric functions as a polynomial ring in countably many variables.
2.3.
For any integer $m$, define the derivations
$\mathsf{D}_{m}\colon\Lambda\to\Lambda\,,\qquad\mathsf{D}_{m}=\sum_{k\geq 0}k^{m}\frac{\partial}{\partial\mathsf{\mathbf{p}}_{k}}.$ (2.1)
In [DFO25b] we considered a derivation $\mathsf{D}_{s}$, where $s$ is a formal variable. There it was showed to be a quasi-isometry, property that allowed us to give recursive formulas to compute tensor products and plethysms of Schur functions. Here we are considering the evaluation of $\mathsf{D}_{s}$ at $s=m$.
Next, we want to explain how to explicitly compute $\mathsf{D}_{m}(\mathsf{\mathbf{s}}_{\alpha})$. The following interpretation is proved in [DFO25b, Section 3] (see also [DFO25b, Proposition 2.10]).
For each $(i,j)\in\alpha$ we denote by $\alpha_{(i,j)}$ the Young diagram obtained by $\alpha$ by removing the hook $\operatorname{Hook}_{(i,j)}(\alpha)$. If removing a hook disconnects the diagram, then we move the lower component up-left.
Example 2.1.
If $\alpha=(3,2,2)$, then $\alpha_{(1,2)}=(3,1)$.
Proposition 2.2 ([DFO25b, (3.1)]).
The following equality holds,
$\mathsf{D}_{m}(\mathsf{\mathbf{s}}_{\alpha})=\sum_{(i,j)\in\alpha}(-1)^{\varepsilon(i,j)}\mathsf{\mathbf{s}}_{\alpha_{(i,j)}}\operatorname{h}_{(i,j)}^{m-1},$
where $\varepsilon(i,j)$ is the number of boxes strictly below $(i,j)$.
Example 2.3.
Continuing Example 2.1, we have
$\mathsf{D}_{m}(\mathsf{\mathbf{s}}_{(3,2,2)})=\mathsf{\mathbf{s}}_{(3,2,1)}+\mathsf{\mathbf{s}}_{2,2,2}+(\mathsf{\mathbf{s}}_{(3,2)}-\mathsf{\mathbf{s}}_{(3,1,1)})2^{m-1}-\mathsf{\mathbf{s}}_{3,1}3^{m-1}+\mathsf{\mathbf{s}}_{1,1,1}4^{m-1}+\mathsf{\mathbf{s}}_{(1,1)}5^{m-1}\,.$
Example 2.4.
For symmetric and exterior powers, we have
$\mathsf{D}_{m}(\mathsf{\mathbf{s}}_{n})=\sum_{k=1}^{n}k^{m-1}\mathsf{\mathbf{s}}_{n-k}\qquad\mbox{and}\qquad\mathsf{D}_{m}(\mathsf{\mathbf{s}}_{1^{n}})=\sum_{k=1}^{n}(-1)^{k+1}k^{m-1}\mathsf{\mathbf{s}}_{1^{n-k}}.$
As in (1.3), we put $\mathsf{D}^{\mu}=\mathsf{D}_{\ell}^{\mu_{\ell}}\cdots\mathsf{D}_{1}^{\mu_{1}}$, where $\mu=(1^{\mu_{1}},\dots,\ell^{\mu_{\ell}})$ is a partition, and notice that $\mathsf{D}_{n}\cdot\mathsf{D}_{m}=\mathsf{D}_{m}\cdot\mathsf{D}_{n}$ for every $m$ and $n$.
Example 2.5.
Since $\mathsf{D}_{1}(\mathsf{\mathbf{s}}_{m})=\mathsf{\mathbf{s}}_{m-1}+\mathsf{\mathbf{s}}_{m-2}+\cdots+1$, it is easy to see that, for every $k>0$,
$\mathsf{D}_{1}^{k}(\mathsf{\mathbf{s}}_{m})=\sum_{j=0}^{m-k}\binom{m-1-j}{k-1}\mathsf{\mathbf{s}}_{j}.$
Similarly, it is also easy to see that $\mathsf{D}_{2}^{k}(\mathsf{\mathbf{s}}_{m})=\mathsf{D}_{1}^{2k}(\mathsf{\mathbf{s}}_{m+k})$.
Remark 2.6.
If $|\alpha|=n$, then $\mathsf{D}^{\mu}(\mathsf{\mathbf{s}}_{\alpha})=0$ for every $\mu=(1^{\mu_{1}},\dots,\ell^{\mu_{\ell}})$ such that $\mu_{1}+\cdots+\mu_{l}>n$. (If we express $\mu$ as a Young diagram, then $\mu_{1}+\cdots+\mu_{l}$ is the number of rows of the diagram.)
3. Chern character of Schur bundles
Through this section we assume that $X$ is a connected compact complex manifold. In this case we consider, as graded commutative algebra $A$, the even cohomology ring $\operatorname{H}^{2\ast}(X,\mathbb{Q})$. Notice that since $\operatorname{H}^{2n}(X,\mathbb{Q})\neq 0$ only for finitely many $n$, we have $A=\widehat{A}$.
Let $E$ be a rank $r$ holomorphic vector bundle on $X$. Consider the following homomorphism
$\mathsf{C}^{E}\colon\Lambda\to\operatorname{H}^{2\ast}(X,\mathbb{Q}),\qquad\mathsf{\mathbf{s}}_{\alpha}\mapsto\operatorname{ch}(\mathsf{\mathbf{S}}^{\alpha}E).$ (3.1)
It is easy to check that it is a characteristic class as in Definition 1.1. Notice that in this case we have full additivity, not just same-degree additivity (cf. Remark 1.2).
We also point out that
$\mathsf{C}^{E}_{0}(\mathsf{\mathbf{s}}_{\alpha})=\operatorname{rk}(\mathsf{\mathbf{S}}^{\alpha}E)=\prod_{(i,j)\in\alpha}\frac{r+i-j}{\operatorname{h}_{(i,j)}(\alpha)},$
where the last equality is the semi-standard tableaux hook length formula.
Remark 3.1.
By the splitting principle we have $\mathsf{C}_{0}^{E}(\mathsf{\mathbf{p}}_{n})=r$ for every $n\geq 1$. This is the only point where the splitting principle is needed.
As a direct application of Corollary 1.4, we get the following formula for the Chern character of Schur bundles.
Theorem 3.2.
With notations as above, we have
$\operatorname{ch}_{n}(\mathsf{\mathbf{S}}^{\alpha}E)=\sum_{\nu\vdash n}\operatorname{rk}(\mathsf{D}^{\nu}(\mathsf{\mathbf{s}}_{\alpha}))\,\operatorname{ch}^{(\nu)}(E).$
Let us see some consequences and explicit examples.
Corollary 3.3.
With notations as above,
$\operatorname{ch}_{1}(\mathsf{\mathbf{S}}^{\alpha}E)=|\alpha|\frac{r_{\alpha}}{r}\operatorname{ch}_{1}(E),$
where $r_{\alpha}=\operatorname{rk}(\mathsf{\mathbf{S}}^{\alpha}E)$.
Proof.
From Theorem 3.2 we obtain that
$\operatorname{ch}_{1}(\mathsf{\mathbf{S}}^{\alpha}E)=\operatorname{rk}(\mathsf{D}_{1}(\mathsf{\mathbf{s}}_{\alpha}))\operatorname{ch}_{1}(E).$
As
$\mathsf{D}_{1}=\sum_{k\geq 0}\frac{\partial}{\partial\mathsf{\mathbf{p}}_{k}}$
and $\operatorname{rk}(\mathsf{\mathbf{p}}_{k})=r$ for every $k$, then
$\operatorname{rk}(\mathsf{D}_{1}(\mathsf{\mathbf{s}}_{\alpha}))=\frac{1}{r}\operatorname{rk}\left(\sum_{k}\mathsf{\mathbf{p}}_{k}\frac{\partial}{\partial\mathsf{\mathbf{p}}_{k}}(\mathsf{\mathbf{s}}_{\alpha})\right).$
However, $\sum_{k}\mathsf{\mathbf{p}}_{k}\frac{\partial}{\partial\mathsf{\mathbf{p}}_{k}}$ acts on homogenous elements via multiplication by their degree (enough to check on a homogeneous basis, like for example the monomials in power sums, where it is evident), thus
$\operatorname{ch}_{1}(\mathsf{\mathbf{S}}^{\alpha}E)=|\alpha|\frac{r_{\alpha}}{r}\operatorname{ch}_{1}(E).$
Remark 3.4.
Corollary 3.3 gives an alternative proof of the main result of [Rub13].
We wish to point out the efficiency of Theorem 3.2 by explicitly computing some examples.
Example 3.5 (Schur bundles of size $2$).
Let us compute $\operatorname{ch}(\mathsf{\mathbf{S}}^{\alpha}E)$ when $|\alpha|=2$. We have only two cases, namely the symmetric power $\mathsf{\mathbf{S}}^{2}E$ and the exterior power $\bigwedge^{2}E$, corresponding to $\alpha=(1,1)$.
$\alpha=(2)$:
it is easy to see that the only non-zero derivatives of $\mathsf{\mathbf{s}}_{2}$ are
$\mathsf{D}_{a}(\mathsf{\mathbf{s}}_{2})=\mathsf{\mathbf{s}}_{1}+2^{a-1}\qquad\mbox{ and }\qquad\mathsf{D}_{b}\mathsf{D}_{a}(\mathsf{\mathbf{s}}_{2})=1.$
Then the first values of $\operatorname{ch}_{n}(\mathsf{\mathbf{S}}^{2}E)$ are:
$\operatorname{rk}(\mathsf{\mathbf{S}}^{2}E)=$ $\,\frac{r(r+1)}{2}$
$\operatorname{ch}_{1}(\mathsf{\mathbf{S}}^{2}E)=$ $\,(r+1)\operatorname{ch}_{1}(E)$
$\operatorname{ch}_{2}(\mathsf{\mathbf{S}}^{2}E)=$ $\,(r+2)\operatorname{ch}_{2}(E)+\frac{1}{2}\operatorname{ch}_{1}(E)^{2}$
$\operatorname{ch}_{3}(\mathsf{\mathbf{S}}^{2}E)=$ $\,(r+4)\operatorname{ch}_{3}(E)+\operatorname{ch}_{2}(E)\operatorname{ch}_{1}(E)$
$\operatorname{ch}_{4}(\mathsf{\mathbf{S}}^{2}E)=$ $\,(r+8)\operatorname{ch}_{4}(E)+\operatorname{ch}_{3}(E)\operatorname{ch}_{1}(E)+\frac{1}{2}\operatorname{ch}_{2}(E)^{2}.$
$\alpha=(1,1)$:
it is easy to see that the only non-zero derivatives of $\mathsf{\mathbf{s}}_{(1,1)}$ are
$\mathsf{D}_{a}(\mathsf{\mathbf{s}}_{(1,1)})=\mathsf{\mathbf{s}}_{1}-2^{a-1}\qquad\mbox{ and }\qquad\mathsf{D}_{b}\mathsf{D}_{a}(\mathsf{\mathbf{s}}_{(1,1)})=1.$
Then the first values of $\operatorname{ch}_{n}(\mathsf{\mathbf{S}}^{2}E)$ are:
$\operatorname{rk}(\bigwedge^{2}E)=$ $\,\frac{r(r-1)}{2}$
$\operatorname{ch}_{1}(\bigwedge^{2}E)=$ $\,(r-1)\operatorname{ch}_{1}(E)$
$\operatorname{ch}_{2}(\bigwedge^{2}E)=$ $\,(r-2)\operatorname{ch}_{2}(E)+\frac{1}{2}\operatorname{ch}_{1}(E)^{2}$
$\operatorname{ch}_{3}(\bigwedge^{2}E)=$ $\,(r-4)\operatorname{ch}_{3}(E)+\operatorname{ch}_{2}(E)\operatorname{ch}_{1}(E)$
$\operatorname{ch}_{4}(\bigwedge^{2}E)=$ $\,(r-8)\operatorname{ch}_{4}(E)+\operatorname{ch}_{3}(E)\operatorname{ch}_{1}(E)+\frac{1}{2}\operatorname{ch}_{2}(E)^{2}.$
Example 3.6 (Schur bundles of size $3$).
In this case we have three cases, namely $\alpha=(3),(2,1),(1,1,1)$.
$\bullet\;\alpha=(3)$:
It is easy to compute
$\mathsf{D}_{a}(\mathsf{\mathbf{s}}_{3})=\mathsf{\mathbf{s}}_{2}+\mathsf{\mathbf{s}}_{1}2^{a-1}+3^{a-1},\quad\mathsf{D}_{b}\mathsf{D}_{a}(\mathsf{\mathbf{s}}_{3})=\mathsf{\mathbf{s}}_{1}+2^{a-1}+2^{b-1}\quad\mbox{and}\quad\mathsf{D}_{c}\mathsf{D}_{b}\mathsf{D}_{a}(\mathsf{\mathbf{s}}_{3})=1,$
from which we deduce the first values of $\operatorname{ch}_{n}(\mathsf{\mathbf{S}}^{3}E)$:
$\operatorname{rk}(\mathsf{\mathbf{S}}^{3}E)=$ $\,\binom{r+2}{3}$
$\operatorname{ch}_{1}(\mathsf{\mathbf{S}}^{3}E)=$ $\,\binom{r+2}{2}\operatorname{ch}_{1}(E)$
$\operatorname{ch}_{2}(\mathsf{\mathbf{S}}^{3}E)=$ $\,\binom{r+3}{2}\operatorname{ch}_{2}(E)+\frac{r+2}{2}\operatorname{ch}_{1}(E)^{2}$
$\operatorname{ch}_{3}(\mathsf{\mathbf{S}}^{3}E)=$ $\,\frac{(r+3)(r+6)}{2}\operatorname{ch}_{3}(E)+(r+3)\operatorname{ch}_{2}(E)\operatorname{ch}_{1}(E)+\frac{1}{6}\operatorname{ch}_{1}(E)^{3}$
$\operatorname{ch}_{4}(\mathsf{\mathbf{S}}^{3}E)=$ $\,\frac{r^{2}+17r+54}{2}\operatorname{ch}_{4}(E)+(r+5)\operatorname{ch}_{3}(E)\operatorname{ch}_{1}(E)+\frac{r+4}{2}\operatorname{ch}_{2}(E)^{2}+\frac{1}{2}\operatorname{ch}_{2}(E)\operatorname{ch}_{1}(E)^{2}.$
$\bullet\;\alpha=(2,1)$:
It is easy to see that
$\mathsf{D}_{a}(\mathsf{\mathbf{s}}_{(2,1)})=\mathsf{\mathbf{s}}_{2}+\mathsf{\mathbf{s}}_{(1,1)}-3^{a-1},\quad\mathsf{D}_{b}\mathsf{D}_{a}(\mathsf{\mathbf{s}}_{(2,1)})=2\mathsf{\mathbf{s}}_{1}\qquad\mbox{and}\qquad\mathsf{D}_{c}\mathsf{D}_{b}\mathsf{D}_{a}(\mathsf{\mathbf{s}}_{(2,1)})=2,$
from which we deduce the first values of $\operatorname{ch}_{n}(\mathsf{\mathbf{S}}^{(2,1)}E)$:
$\operatorname{rk}(\mathsf{\mathbf{S}}^{(2,1)}E)=$ $\,\frac{r(r^{2}-1)}{3}$
$\operatorname{ch}_{1}(\mathsf{\mathbf{S}}^{(2,1)}E)=$ $\,(r^{2}-1)\operatorname{ch}_{1}(E)$
$\operatorname{ch}_{2}(\mathsf{\mathbf{S}}^{(2,1)}E)=$ $\,(r^{2}-3)\operatorname{ch}_{2}(E)+r\operatorname{ch}_{1}(E)^{2}$
$\operatorname{ch}_{3}(\mathsf{\mathbf{S}}^{(2,1)}E)=$ $\,(r^{2}-9)\operatorname{ch}_{3}(E)+2r\operatorname{ch}_{2}(E)\operatorname{ch}_{1}(E)+\frac{\operatorname{ch}_{1}(E)^{3}}{3}$
$\operatorname{ch}_{4}(\mathsf{\mathbf{S}}^{(2,1)}E)=$ $\,(r^{2}-27)\operatorname{ch}_{4}(E)+2r\operatorname{ch}_{3}(E)\operatorname{ch}_{1}(E)+r\operatorname{ch}_{2}(E)^{2}+\operatorname{ch}_{2}(E)\operatorname{ch}_{1}(E)^{2}.$
$\bullet\;\alpha=(1,1,1)$:
It is easy to compute
$\mathsf{D}_{a}(\mathsf{\mathbf{s}}_{(1,1,1)})=\mathsf{\mathbf{s}}_{(1,1)}-\mathsf{\mathbf{s}}_{1}2^{a-1}+3^{a-1},\quad\mathsf{D}_{b}\mathsf{D}_{a}(\mathsf{\mathbf{s}}_{3})=\mathsf{\mathbf{s}}_{1}-2^{a-1}-2^{b-1}\quad\mbox{and}\quad\mathsf{D}_{c}\mathsf{D}_{b}\mathsf{D}_{a}(\mathsf{\mathbf{s}}_{3})=1,$
from which we deduce the first values of $\operatorname{ch}_{n}(\bigwedge^{3}E)$:
$\operatorname{rk}(\bigwedge^{3}E)=$ $\,\binom{r}{3}$
$\operatorname{ch}_{1}(\bigwedge^{3}E)=$ $\,\binom{r-1}{2}\operatorname{ch}_{1}(E)$
$\operatorname{ch}_{2}(\bigwedge^{3}E)=$ $\,\binom{r-2}{2}\operatorname{ch}_{2}(E)+\frac{r-2}{2}\operatorname{ch}_{1}(E)^{2}$
$\operatorname{ch}_{3}(\bigwedge^{3}E)=$ $\,\frac{(r-3)(r-6)}{2}\operatorname{ch}_{3}(E)+(r-3)\operatorname{ch}_{2}(E)\operatorname{ch}_{1}(E)+\frac{1}{6}\operatorname{ch}_{1}(E)^{3}$
$\operatorname{ch}_{4}(\bigwedge^{3}E)=$ $\,\frac{r^{2}-17r+54}{2}\operatorname{ch}_{4}(E)+(r-5)\operatorname{ch}_{3}(E)\operatorname{ch}_{1}(E)+\frac{r-4}{2}\operatorname{ch}_{2}(E)^{2}+\frac{1}{2}\operatorname{ch}_{2}(E)\operatorname{ch}_{1}(E)^{2}.$
3.1. Chern character of symmetric and exterior powers
It will be useful to introduce the following notations. If $\nu=(1^{\nu_{1}},\dots,\ell^{\nu_{\ell}})$ is a partition, then we denote by
$\mathbf{a}_{\nu}=(\overbrace{\ell,\dots,\ell}^{\nu_{\ell}},\dots,\overbrace{1,\dots,1}^{\nu_{1}})$
the associated Young diagram. We denote by $a_{i}$ the entries of $\mathbf{a}_{\nu}$. Accordingly we write
$\mathsf{D}^{\nu}=\mathsf{D}_{\mathbf{a}_{\nu}}=\overbrace{\mathsf{D}_{\ell}\circ\cdots\circ\mathsf{D}_{\ell}}^{\nu_{\ell}}\circ\cdots\circ\overbrace{\mathsf{D}_{1}\circ\cdots\circ\mathsf{D}_{1}}^{\nu_{1}}.$
Let us set
$h=h(\nu)=\nu_{1}+\cdots\nu_{\ell}$
to be the height of the Young diagram $\mathbf{a}_{\nu}$, i.e. the number of rows. If $\mathbf{b}=(b_{1},\dots,b_{h})$ is $h$-tuple of integers, we put
$\overleftarrow{\mathbf{b}}=(b_{1}-1,\dots,b_{h}-1).$
Finally, for every $n,j\geq 0$, let us denote by $\mathsf{\mathbf{E}}(n,j)$ the corresponding Eulerian number. Recall the following important property of Eulerian numbers:
$k^{a}=\sum_{b=1}^{a+1}\mathsf{\mathbf{E}}(a,b-1)\binom{k+b-1}{a}$. (3.2)
Remark 3.7.
Let us notice that $\mathsf{\mathbf{E}}(n,k)=0$ for all $k\geq n$, unless $\mathsf{\mathbf{E}}(0,0)$ which is set equal to 11 by definition. Formula (3.2) is usually written with $b$ ranging from 11 to $a$, but would leave the case $a=0$ out, which we need in the proof of Proposition 3.8.
If $\mathbf{b}$ and $\mathbf{c}$ are two $h$-tuples of natural numbers, then we put
$\mathsf{\mathbf{E}}(\mathbf{c},\mathbf{b})=\prod_{i=1}^{h}\mathsf{\mathbf{E}}(c_{i},b_{i}).$
Proposition 3.8.
With notations as above, we have
$\operatorname{ch}_{n}(\mathsf{\mathbf{S}}^{m}E)=\sum_{\nu\vdash n}\sum_{\mathbf{b}\leq\mathbf{a}_{\nu}}\mathsf{\mathbf{E}}(\overleftarrow{\mathbf{a}_{\nu}},\overleftarrow{\mathbf{b}})\binom{r+m+|\mathbf{b}|-\nu_{1}-1}{r+|\mathbf{a}_{\nu}|-1}\,\operatorname{ch}^{(\nu)}(E)$ (3.3)
and
$\operatorname{ch}_{n}\left(\bigwedge^{m}E\right)=\sum_{\nu\vdash n}\sum_{\mathbf{b}\leq\mathbf{a}_{\nu}}\mathsf{\mathbf{E}}(\overleftarrow{\mathbf{a}_{\nu}},\overleftarrow{\mathbf{b}})\binom{r-|\mathbf{a}_{\nu}|}{m-|\mathbf{a}_{\nu}|+|\mathbf{b}|-\nu_{1}}\,\operatorname{ch}^{(\nu)}(E),$ (3.4)
where the sums run over all $\mathbf{b}=(b_{1},\dots,b_{h(\nu)})$ such that $0<b_{i}\leq a_{i}$ for every $i$, and $|\mathbf{b}|=\sum_{i}b_{i}$.
Remark 3.9.
We can be more precise about the bounds $0<b_{i}\leq a_{i}$. In fact, if $a_{i}>1$, then we have $0<b_{i}\leq a_{i}-1$, as it will be clear from the proof. The extremal case $b_{i}=a_{i}$ only appears when $a_{i}=1$, i.e. $a_{i}-1=0$, see also Remark 3.7.
In the proof we will need the following well-known binomial identities.
Lemma 3.10.
For any non-negative integers $A$, $R$ and $N$,
$\sum_{i=0}^{N}\binom{A+i}{A}\binom{R+N-i}{R}=\binom{R+N+A+1}{R+A+1}$ (3.5)
and
$\sum_{i=0}^{N}\binom{A+i}{A}\binom{R}{N-i}=\binom{R-A-1}{N}$. (3.6)
Proof of Proposition 3.8 (symmetric power case).
It is enough to prove that
$\operatorname{rk}(\mathsf{D}_{\mathbf{a}_{\nu}}(\mathsf{\mathbf{s}}_{m}))=\sum_{\mathbf{b}\leq\mathbf{a}_{\nu}}\mathsf{\mathbf{E}}(\overleftarrow{\mathbf{a}_{\nu}},\overleftarrow{\mathbf{b}})\binom{r+m+|\mathbf{b}|-\nu_{1}-1}{r+|\mathbf{a}_{\nu}|-1}$. (3.7)
We proceed by induction on $h=h(\nu)$. As base of the induction, let us prove the $\mathbf{a}_{\nu}=(a)$. By definition and (3.2),
$\operatorname{rk}(\mathsf{D}_{a}(\mathsf{\mathbf{s}}_{m}))=\sum_{k=1}^{m}k^{a-1}\binom{r+m-k-1}{r-1}=\sum_{b=1}^{a}\mathsf{\mathbf{E}}(a-1,b-1)\sum_{k=1}^{m}\binom{k+b-1}{a-1}\binom{r+m-k-1}{r-1}$.
The sum on the right can be computed using Lemma 3.10, but it is important to distinguish two cases. First, if $a=1$, then the sum reduces to the well-known equality $\sum_{k=1}^{m}\binom{r+m-k-1}{r-1}=\binom{r+m-1}{r}$, that agrees with (3.7), since in this case only $b=1$ is allowed and $\nu_{1}=1$. (Notice that in this case, contrary to the next one, the sum starts with $k=1$; this is the reason why we have an extra 1-1 in the numerator of the binomial coefficient.) If $a\geq 2$, then we have non-trivial summands only for $a-b\leq k\leq m$ so that we can apply Lemma 3.10 with A=a1A=a-1, R=r1R=r-1 and N=ma+bN=m-a+b, yielding
$\operatorname{rk}(\mathsf{D}_{a}(\mathsf{\mathbf{s}}_{m}))=\sum_{b=1}^{a-1}\mathsf{\mathbf{E}}(a-1,b-1)\binom{r+m+b-1}{r+a-1}$.
Since $\nu_{1}=0$ in this case, this agrees again with (3.7).
For the inductive step, let us first assume that there exists an index $i$ such that $a_{i}>1$. Without loos of generality, we can assume that it is $a_{1}>1$. In fact, since $\mathsf{D}_{a_{i}}$ and $\mathsf{D}_{a_{j}}$ commute we can re-order the indices. Let us then write $\mathbf{a}_{\nu}=(a_{1},\dots,a_{h})$, and let us set $\mathbf{a}^{\prime}=(a_{2},\dots,a_{h})$. By definition we have
$\operatorname{rk}(\mathsf{D}_{\mathbf{a}}(\mathsf{\mathbf{s}}_{m}))=\sum_{k=1}^{m}k^{a_{1}-1}\operatorname{rk}(\mathsf{D}_{\mathbf{a}^{\prime}}(\mathsf{\mathbf{s}}_{m-k})).$
By (3.2) and the inductive hypothesis, we have
$\operatorname{rk}(\mathsf{D}_{\mathbf{a}}(\mathsf{\mathbf{s}}_{m}))=\sum_{b_{1}=1}^{a_{1}-1}\sum_{\mathbf{b}^{\prime}\leq\mathbf{a}^{\prime}}\mathsf{\mathbf{E}}(\overleftarrow{\mathbf{a}^{\prime}},\overleftarrow{\mathbf{b}^{\prime}})\mathsf{\mathbf{E}}(a_{1}-1,b_{1}-1)\sum_{k=1}^{m}\binom{k+b_{1}-1}{a_{1}-1}\binom{r+m-k+|\mathbf{b}^{\prime}|-\nu_{1}-1}{r+|\mathbf{a}^{\prime}|-1}$,
where we write $\mathbf{b}^{\prime}=(b_{2},\dots,b_{h})$. As before, the last sum has non-trivial summands only for $a_{1}-b_{1}\leq k\leq m-|\mathbf{a}^{\prime}|+|\mathbf{b}^{\prime}|-\nu_{1}$ and we can apply Lemma 3.10 with $A=a_{1}-1$, $R=r+|\mathbf{a}^{\prime}|-1$ and $N=m-|\mathbf{a}|+|\mathbf{b}|-\nu_{1}$ (here $\mathbf{b}=(b_{1},\dots,b_{h})$), resulting in
$\sum_{k=1}^{m}\binom{k+b_{1}-1}{a_{1}-1}\binom{r+m-k+|\mathbf{b}^{\prime}|-\nu_{1}-1}{r+|\mathbf{a}^{\prime}|-1}=\binom{r+m+|\mathbf{b}|-\nu_{1}-1}{r+|\mathbf{a}|-1}$,
which is again (3.7).
Finally, if $a_{1}=\cdots=a_{h}=1$, then we can either proceed as above (using the formula for $\operatorname{rk}\mathsf{D}_{1}(\mathsf{\mathbf{s}}_{m})$ above) or we can directly compute $\operatorname{rk}\mathsf{D}_{1}^{h}(\mathsf{\mathbf{s}}_{m})$ thanks to Example 2.5. In both cases we get
$\operatorname{rk}\mathsf{D}_{1}^{h}(\mathsf{\mathbf{s}}_{m})=\binom{r+m-1}{r+h-1}$,
which again agrees with (3.7), thus concluding the proof. ∎
Proof of Proposition 3.8 (exterior power case).
The proof is analogous to the previous one, so we will only sketch it.
First of all, let us deal with the case $\mathbf{a}_{\nu}=(a)$. As before, we have different behavior if $a=1$ or $a>1$. Since
$\operatorname{rk}\mathsf{D}_{a}\left(\bigwedge^{m}E\right)=\sum_{k01}^{m}(-1)^{k+1}k^{a-1}\binom{r}{m-k}=\left\{\begin{array}[]{ll}\binom{r-1}{m-1}&\mbox{if }a=1\\ \sum_{b=1}^{a-1}\mathsf{\mathbf{E}}(a-1,b-1)\binom{r-a}{m-a-b}&\mbox{if }a>1\,.\end{array}\right$.
Let now $\mathbf{a}_{\nu}=(a_{1},\dots,a_{h})$ and assume that for at least one index $i$ we have $a_{i}>1$. Without loss of generality, it is $a_{1}>1$. If $\mathbf{a}^{\prime}=(a_{2},\dots,a_{h})$, then
$\operatorname{rk}\mathsf{D}_{\mathbf{a}_{\nu}}\left(\bigwedge^{m}E\right)=$ $\,\sum_{k=1}^{m}(-1)^{k+1}k^{a_{1}-1}\operatorname{rk}\mathsf{D}_{\mathbf{a}^{\prime}}\left(\bigwedge^{m-k}E\right)$
$=$ $\,\sum_{b_{1}=1}^{a_{1}-1}\sum_{\mathbf{b}^{\prime}\leq\mathbf{a}^{\prime}}\mathsf{\mathbf{E}}(\overleftarrow{\mathbf{a}^{\prime}},\overleftarrow{\mathbf{b}^{\prime}})\mathsf{\mathbf{E}}(a_{1}-1,b_{1}-1)\sum_{k=1}^{m}(-1)^{k+1}\binom{k+b_{1}-1}{a_{1}-1}\binom{r-|\mathbf{a}^{\prime}|}{m-|\mathbf{a}^{\prime}|+|\mathbf{b}^{\prime}|-\nu_{1}}$
$=$ $\,\sum_{\mathbf{b}\leq\mathbf{a}_{\nu}}\mathsf{\mathbf{E}}(\overleftarrow{\mathbf{a}_{\nu}},\overleftarrow{\mathbf{b}})\binom{r-|\mathbf{a}_{\nu}|}{m-|\mathbf{a}_{\nu}|+|\mathbf{b}|-\nu_{1}}$,
where the last equality follows from (3.6).
Finally, the case when $a_{1}=\cdots=a_{h}=1$ is done similarly and it yields
$\operatorname{rk}\mathsf{D}_{1}^{h}\left(\bigwedge^{m}E\right)=\binom{r-1}{r-h}$,
thus concluding the proof. ∎
Remark 3.11.
In [Svr93], Svrtan obtains a similar formulas involving Stirling numbers of the second kind. It is known that Eulerian numbers and Stirling numbers of the second kind are related, so our formula could have deduced from his. Nevertheless, notice that our proof is completely different from his.
4. Twisted Chern character of Schur bundles
Let $\mathsf{C}^{E}$ be the characteristic class defined in (3.1). We can twist it in the following way,
$\mathsf{C}^{\mathbb{P}E}\colon\Lambda\to\operatorname{H}^{2\ast}(X,\mathbb{Q}),\qquad\mathsf{\mathbf{s}}_{\alpha}\mapsto k(\mathsf{\mathbf{S}}^{\alpha}E):=\exp\left(-\frac{|\alpha|}{r}\operatorname{ch}_{1}(E)\right)\operatorname{ch}(\mathsf{\mathbf{S}}^{\alpha}E).$
It is easy to check that this is again a characteristic class. Notice that in this case $\mathsf{C}^{\mathbb{P}E}$ is only same-degree additive. The name and the notation are motivated by [Mar20], where these classes are used as Chern classes of projective bundles. More precisely, for any (twisted) vector bundle $F$, Markman defines
$k(F)=\exp\left(-\frac{\operatorname{ch}_{1}(F)}{\operatorname{rk}(F)}\right)\operatorname{ch}(F).$
Our definition is equivalent to the above one thanks to Corollary 3.3. The first three values of $k(F)$ in terms of Chern classes are
$k_{0}(F)=\operatorname{rk}(F),\quad k_{1}(F)=0,\quad k_{2}(F)=-\frac{\operatorname{ch}_{1}(F)^{2}-2\operatorname{rk}(F)\operatorname{ch}_{2}(F)}{2\operatorname{rk}(F)}$
and
$k_{3}(F)=\frac{\operatorname{ch}_{1}(F)^{3}-3\operatorname{rk}(F)\operatorname{ch}_{2}(F)\operatorname{ch}_{1}(F)+3\operatorname{rk}(F)^{2}\operatorname{ch}_{3}(F)}{3\operatorname{rk}(F)^{2}}.$
Notice that $2\operatorname{rk}(F)\,k_{2}(F)=\Delta(F)$ is the discriminant of $F$. As a direct application of Corollary 1.4, we get the following formula for the class $k$ of Schur bundles.
Theorem 4.1.
With notations as above,
$k_{n}(\mathsf{\mathbf{S}}^{\alpha}E)=\sum_{\nu\vdash n}\operatorname{rk}(\mathsf{D}^{\nu}(\mathsf{\mathbf{s}}_{\alpha}))\,k^{(\nu)}(E).$
Example 4.2.
The first values of $k_{n}(\mathsf{\mathbf{S}}^{\alpha}E)$ are easy to compute. In fact, since $k_{1}(E)=0$, we get
$k_{2}(\mathsf{\mathbf{S}}^{\alpha}E)=$ $\,\operatorname{rk}(\mathsf{D}_{2}(\mathsf{\mathbf{s}}_{\alpha}))\,k_{2}(E)\,;$
$k_{3}(\mathsf{\mathbf{S}}^{\alpha}E)=$ $\,\operatorname{rk}(\mathsf{D}_{3}(\mathsf{\mathbf{s}}_{\alpha}))\,k_{3}(E)\,;$
$k_{4}(\mathsf{\mathbf{S}}^{\alpha}E)=$ $\,\operatorname{rk}(\mathsf{D}_{4}(\mathsf{\mathbf{s}}_{\alpha}))\,k_{4}(E)+\frac{1}{2}\operatorname{rk}(\mathsf{D}_{2}^{2}(\mathsf{\mathbf{s}}_{\alpha}))k_{2}^{2}(E)\,$,
so that, for instance,
$k_{2}(\mathsf{\mathbf{S}}^{2}E)=$ $\,(r+2)\,k_{2}(E)\,;$
$k_{3}(\mathsf{\mathbf{S}}^{2}E)=$ $\,(r+4)\,k_{3}(E)\,;$
$k_{4}(\mathsf{\mathbf{S}}^{2}E)=$ $\,(r+8)\,k_{4}(E)+\frac{1}{2}k_{2}^{2}(E)\,.$
Remark 4.3.
It is worth noticing that the class $k$ is related to the logarithmic Chern character as defined in [Dre95]. If we denote by $\mathsf{L}(E)$ the latter, then for any vector bundle $F$,
$\operatorname{Log}(k(F))=\mathsf{L}(E)-\left(\frac{\operatorname{ch}_{1}(F)}{\operatorname{rk}(F)}\right),$ (4.1)
from which it follows, for example, that
$\mathsf{L}_{2}(F)=\frac{k_{2}(F)}{\operatorname{rk}(F)},\quad\mathsf{L}_{3}(F)=\frac{k_{3}(F)}{\operatorname{rk}(F)}\quad\mbox{and}\quad\mathsf{L}_{4}(F)=\frac{k_{4}(F)}{\operatorname{rk}(F)}-\frac{k_{2}(F)^{2}}{\operatorname{rk}(F)^{2}}.$
In [DFO25a] we proved that the classes $\mathsf{L}_{2}(F)$ and $\mathsf{L}_{3}(F)$ are eigen-classes with respect to taking the Schur functors of a bundle. Moreover, we also remarked that the same did not hold for higher-degree classes and speculated that similar formulas would work for the projection of the classes $\mathsf{L}_{n}(\mathsf{\mathbf{S}}^{\alpha}E)$ onto the $\mathsf{L}_{n}(E)$-component.
Theorem 4.1 solves this problem in the best possible way. In fact, as it is clear from Example 4.2, both $k_{2}$ and $k_{3}$ are eigen-classes, thus recovering the result for $\mathsf{L_{2}}$ and $\mathsf{L_{3}}$ in [DFO25a]. Moreover, Theorem 4.1 provides a formula not just for the $\mathsf{L}_{n}(E)$-component, but for all components (up to re-ordering according to (4.1)).